hello may i have help for this question?
The perimeter of a rectangular field is 366 meters. The longest
side of the field is 35 meters longer than the shorter side. What
is the width of the field?

Answer by ewatrrr(24785)

 
Hi,

P = 2L + 2w = 366
L + w = 183 |L = w + 35
(w+35) + w = 183
2w = 148
w = 74m